# Black-box optimization and quantum annealing for filtering out mislabeled training instances

### v1

@@ Introduction
Several studies have examined the benefits of data cleansing when faced with contaminated datasets. For example, removing outliers or mislabeled instances before training has improved classification accuracy [5, 6]. Moreover, the benefits of eliminating harmful training instances often surpass those of hyperparameter optimization [7]. Previous work has used outlier detection methods, noise removal techniques, or ensembles of different algorithms to eliminate mislabeled instances [8]. Another approach involves removing instances that are mostly misclassified by a set of classifiers[5, 7]. Furthermore, some studies indicate that noise in class labels has a greater negative impact on the performance of the induced model than noise in input attributes [9, 10]. All of these studies indicate that the removal of detrimental instances significantly improves the quality of the learned model. However, these existing approaches have shown limited success in removing noise from contaminated training datasets due to the inherent challenges of the problem. Specifically, removing mislabeled instances is a combinatorial optimization problem. The number of possible data point combinations increases exponentially with the size of the dataset, making brute-force methods impractical.
To address these challenges, this study proposes an approach to noise removal that combines surrogate model-based black-box optimization (BBO) [11, 12, 13, 14] with postprocessing [15] and quantum annealing [16, 17, 18]. Quantum annealing is particularly well suited to solve combinatorial optimization problems [19, 20], due to its demonstrated ability to efficiently sample diverse and high-quality solutions [21, 22]. By leveraging these strengths, the proposed method identifies and effectively removes mislabeled instances from potentially contaminated training datasets.
Our approach utilizes (1) a concrete validation loss metric for selecting data points under the assumption of a clean validation dataset, (2) a sequential training subset selection strategy utilizing BBO with a surrogate model estimating validation loss, and (3) the unique capabilities of quantum annealers to sample diverse and high-quality solutions efficiently. Together, these components address the challenges of noise removal in a structured and scalable manner.
@@ Methods
To evaluate the effectiveness of the proposed algorithm in removing mislabeled instances, we designed a classification task called the noisy majority bit task. This task aims to predict the majority bit in a given binary input vector. The dataset for this task is partitioned into distinct, non-overlapping training, validation, and test sets to ensure an unbiased evaluation. For instance, the correct target label is 00 for the binary input vector $[0,0,0,1,0,1,0,0,1]$, whereas $1$ for $[1,1,0,1,0,1,1,0,0]$. Under ideal conditions, where all training instances are accurately labeled according to the majority bit, this task would be straightforward. However, the actual training set includes noisy instances where the labels are flipped to the minority bit, introducing mislabeled data. This contamination makes the noisy majority bit task non-trivial, providing a controlled and challenging framework for evaluating the proposed algorithm’s effectiveness in noise removal.
@@
Although validation loss can be used as a proxy for the generalization performance of a trained model, it is computationally prohibitive to evaluate the validation loss for all possible subsets of the original training dataset. To address this issue, we employed surrogate model-based BBO [11, 12, 13, 14] with postprocessing [15] to efficiently identify training subsets with low validation error in an iterative manner. It is important to note that while surrogate model-based BBO with postprocessing is not a novel approach, its application to noise removal and the incorporation of diverse samples obtained from quantum annealer during the sample acceptance process are remarkable contribution of this work. The following part explains the detailed procedure of this algorithm and its application to removal of mislabeled training instances. The pseudo code of the algorithm is shown in Algorithm 1.
@@
1:$D_{\mathrm{train}}\equiv\{(\bm{x}_{1},t_{1}),\ldots,(\bm{x}_{n},t_{n})\},D_{\mathrm{valid}}$, QUBO-based sampler, Number of samples drawn from the sampler $M$, Total number of steps $N$; Number of initialization steps: $N_{0}$, Regression model $R$, Supervised-learning model $P$, Monotonically increasing function $g$
2:$\bm{q}_{k^{*}}$ with the lowest validation loss
3:for $k=1$ to $N$ do
@@
to estimate the validation loss of the the model trained by the filtered training subset $D_{\mathrm{train}}(\bm{q})$. In the surrogate model-based BBO, the surrogate model plays two roles: one is to approximate the unknown cost function with a quadratic binary form and the other is to generate binary samples with sufficiently low cost values. Here, the surrogate model $f_{\tilde{\bm{\alpha}}}(\bm{q})$ is parameterized by the parameter vector $\tilde{\bm{\alpha}}\in\mathbb{R}^{p}$ given in the form of Eq. (4) where $p=1+n+n(n-1)/2$. The elements in this parameter vector $\tilde{\bm{\alpha}}$ are also found in the constant-shifted quadratic unconstrained binary optimization (QUBO) form in Eq. (2). Eq. (2) is composed of the first term $\alpha_{0}$ which represents a constant shift and the second quadratic term which represents linear interaction by its diagonal elements and pair-wise interaction by its off-diagonal elements. In short, the unknown true cost function $f(\bm{q})$ over $\bm{q}\in\{0,1\}^{n}$ is approximated by the second order Taylor expansion around $\bm{q}$, which is $f_{\tilde{\bm{\alpha}}}(\bm{q})$ in the form shown in Eq. (3). The quality of this approximation can be improved by adjusting the parameter vector $\tilde{\bm{\alpha}}$.
@@
In the first phase, we randomly samples the $n$-dimensional binary pattern $\bm{q}_{k}\in\{0,1\}^{n}$ sequentially from $k=1$ to $k=N_{0}$, and then for each binary pattern $\bm{q}_{k}$, we compute the validation loss $L_{\mathrm{valid}}(\bm{q}_{k})$ of the model trained by the filtered training subsets $D_{\mathrm{train}}(\bm{q}_{k})$. The validation loss $L_{\mathrm{valid}}(\bm{q}_{k})$ might not be distributed normally, and it is often preferable to transform it by a monotonically increasing function $g$ such as the log function to make it more normally distributed when this value becomes the target of some model and this is the case in our algorithm. Let us define the transformed validation loss as $l_{k}=g(L_{\mathrm{valid}}(\bm{q}_{k}))$. If we define $\tilde{\bm{q}}$ as in Eq. (4), the transformed validation loss $l_{k}$ can be approximated by the surrogate model $f_{\tilde{\bm{\alpha}}}(\bm{q}_{k})$ in the from of Eq. (5), which is linear in $\tilde{\bm{q}}$. Therefore, if we collect the transformed validation losses $\{l_{1},\ldots,l_{N_{0}}\}$ and the corresponding binary patterns $\{\bm{q}_{1},\ldots,\bm{q}_{N_{0}}\}$, we can update the parameter vector $\tilde{\bm{\alpha}}$ to minimize the transformed validation losses $\{l_{1},\ldots,l_{N_{0}}\}$. This optimization can be done in many forms such as the Bayesian linear regression with sparse prior [11, 23, 24], the LASSO regression [25], and the Ridge regression to update the parameter vector $\tilde{\bm{\alpha}}$ depending on the sparsity assumption of the given problem. This first step is necessary to obtain a good initial guess of the parameter vector $\tilde{\bm{\alpha}}$.
@@
In the second phase, we sequentially accept one good sample per step and compute the parameter vector $\tilde{\bm{\alpha}}_{k}$ with sequentially updated input-output pairs $\{(\tilde{\bm{q}}_{j},l_{j})\}_{j=1}^{k-1}$ available at each time step $k\in\{N_{0}+1,N_{0}+2,\ldots,N\}$. The good sample is selected as follows. After drawing $M$ samples $\{\bm{q}_{k}^{(m)}\}_{m=1}^{M}$ from a sampler at step $k$, the one with the lowest energy that is not included in the accepted sample set $S_{k-1}\equiv\{\bm{q}_{1},\ldots,\bm{q}_{k-1}\}$ is selected. If no such sample exists, a random sampling is conducted until a previously unseen sample is found, which is then accepted. The accepted sample $\bm{q}_{k}$ at step $k$ is then used to compute the transformed validation loss $l_{k}$ at step $k$, by transforming the validation loss $L_{\mathrm{valid}}(\bm{q}_{k})$ of the model trained by a filter training subset $D_{\mathrm{train}}(\bm{q}_{k})$. Newly obtained quantities at time step $k$ such as $\bm{q}_{k}\in\{0,1\}^{n}$, $\tilde{\bm{q}}_{k}\in\{0,1\}^{p}$, $l_{k}\in\mathbb{R}$ are stored for the next iteration. At the end of the second phase, the best sample $q_{k^{*}}$ with the lowest transformed validation cost is reported where $k^{*}=\mathrm{argmin}_{k}l_{k}$.
@@
In our experiment, we used the following parameters and settings. $n=128$, $N_{0}=64$, $N=320(=64+256)$, $M=512$, $p=8257=1+128+128\times(128-1)/2$. We used a logistic regression model as a task-specific base model for solving the noisy majority bit task. The reason for using this model is that it is simple, interpretable, and fast to train. As a linear model for estimating $\tilde{\bm{\alpha}}$, we used the ridge regression model instead of the Bayesian model with sparsity prior or the Lasso model, which is usually used in the context of the surrogate model-based BBO to impose the sparsity constraint on the QUBO matrix. The reason for not using sparse-version of the model is that the data is generated in such a way that half of the training data points have noisy labels, and we expect the resulting QUBO matrix can be well approximated without sparsity assumption.
@@ Results
Figure 2: Instance removal patterns acquired by three different samplers over 32 runs. Each cell in the grid represents the inclusion or exclusion status of a training instance for a specific run. For any sampler, a white cell at the $r$-th row and $i$-th column indicates the $i$-th training instance is included in the filtered training set and used for actual training in the $r$-th run. The black cell, on the other hand, shows the same training instance is excluded from the filtered training set and not used for training in the associated run. Seed values are changed from 0 to 31 to see how consistently the proposed algorithm coupled with a certain sampler filters out mislabeled training instances and retains correctly labeled training instances.
@@
A logistic regression model trained on the cleansed training dataset consistently demonstrated improved performance across all 32 runs as shown in Fig. 3. Here, we focus on analyzing the results obtained using the D-Wave QA clique sampler, as similar conclusions are likely applicable to the other samplers. The training errors computed on the optimized training subsets were consistently lower than the training errors computed on the original training sets without noise removal. This indicates that the mislabeled instances were successfully filtered out from the contaminated training dataset. Regardless of successful removal of the mislabeled instances, the training errors computed on the filtered training subsets remained consistently higher than the validation and test errors. This discrepancy arises because some mislabeled instances persist in in the filtered dataset even after noise removal by the proposed algorithm. Please note that the validation error is lowest because it is used as the objective function of the surrogate model-based BBO, and all instances in the validation dataset are noise free. The low test errors suggest that the proposed algorithm effectively removed mislabeled instances in a way to guarantee the generalization of the trained model.
@@
Figure 3: Performance of the trained model on four different datasets. In each run, a model was trained using a filtered training dataset and then evaluated on four different datasets. The model performance was measured in log-loss error metric. Each gray dot represents one of the 32 log-loss errors for a specific dataset. Variations in error values within the same dataset are due to differences in the training instances selected across different runs by our dataset optimization algorithm. The labels Train (all) and Train (optimized) indicate whether all or filtered instances in the training dataset were used for model evaluation, respectively. The whiskers extend to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot.
@@
Further analysis revealed that mislabeled instances with a greater negative impact if not removed are more likely to be eliminated in an input-feature-dependent fashion. Fig. 2 shows that the difference in removal probabilities of fake instances are persistent over different runs and different samplers. In other words, there is a pattern in removal probability of fake instances. Fig. 4 shows the 64 input vectors contained in the fake training instances and their characteristics, such as the summed value of each input vector, together with their removal probabilities computed after 32 runs with the D-Wave QA clique sampler. The relationship between the summed values of input patterns and the removal probabilities of their associated instances is shown in Fig. 5(a). Specifically, the summed input pattern, $s_{i}=\sum_{c=1}^{9}x_{i,c}$, or equivalently the probability of an element of the input vector taking the value of 1, $p_{i}=s_{i}/9$, is shown to influence the removal probability of the input pattern $\bm{x}_{i}\in\{0,1\}^{9}$. To clarify this relationship, Fig 5(b) presents the absolute deviance of summed input pattern from the optimal threshold of 4.5, which is optimal for solving the majority bit task. This absolute deviance defined as $d(\bm{x}_{i})=|9p_{i}-4.5|$ is computed for 64 unique input patterns in the training set $\{\bm{x}_{i}\}_{i=1}^{64}$. Notably, this absolute deviance is inversely related to the entropy of a binary input pattern, defined as $s(\bm{x}_{i})=-p_{i}\log p_{i}-(1-p_{i})\log(1-p_{i})$. Therefore, Fig 5(b) demonstrates that mislabeled input patterns with lower entropy, which correspond to higher absolute deviance, are more likely to be removed. This trend may be attributed to the stronger negative impact of low-entropy patterns with incorrect labels (e.g., $[0,0,0,0,0,0,0,0,0]$ labeled as $1$) on the trained model, compared to high-entropy patterns with incorrect labels (e.g., $[0,1,0,1,0,1,0,1,0]$ labeled as $1$).
@@
Figure 4: Characteristics of the input patterns of the fake training instances and their removal probabilities using the D-Wave QA clique sampler. 1) The 9 dimensional binary input patterns of the fake training instances. 2) The summed values of input patterns, with the horizontal dashed red line marking the optimal decision boundary of 4.5. 3) The absolute deviance (or distance) of summed input patterns from the optimal decision boundary of 4.5. 4) The removal probabilities of incorrectly labeled fake instances over 32 runs with different seed values ranging from 0 to 31. 5) The removal patterns of fake training instances across 32 different runs. This removal pattern is also depicted on the right side of the dashed red line in the D-Wave QA clique result shown in Fig. 2.
@@
Figure 5: Relationship between the removal probabilities and the characteristics of 64 input patterns of the fake training subset. The removal probabilities are compared with (a) the number of ones in each input pattern of the fake training subset and (b) the absolute deviance from the optimal threshold of 4.5 to the summed values of input patterns. In both plots, the whiskers are drawn to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot.
@@
Figure 6: Comparison of three samplers in terms of surrogate losses and types of sequentially obtained samples. The first 64 steps up to the dotted gray line are called the initialization phase, during which samples are obtained randomly. The remaining 256 steps constitute the optimization phase, where samples are obtained using the surrogate model-based BBO with postprocessing. In each optimization step, 512 samples are drawn from a selected sampler. These samples are sequentially checked from the one with the lowest energy to the one with the highest energy to see if they are already contained in the accepted sample set. The first sample not contained in the accepted sample set is accepted and inserted into the accepted sample set. If all 512 samples are already contained in the accepted sample set, a new sample not contained in the acquired sample set is drawn randomly. The trend of the lowest energy up to that step is shown with a blue solid line. A sample type is marked as optimal if the accepted sample has the lowest energy among the 512 samples. A sample type is marked as suboptimal if it is from the 512 samples but does not have the lowest energy. A sample type is marked as random if it is not obtained from the 512 samples and is drawn randomly. In all three cases, all parameters are fixed except for the solver types.

### v2

# Filtering out mislabeled training instances using black-box optimization and quantum annealing

@@ Introduction
Numerous studies have demonstrated that removing noisy or harmful training instances can significantly improve model performance on contaminated datasets. For example, removing outliers or mislabeled instances before training has improved classification accuracy [5, 6]. Moreover, the benefits of eliminating harmful training instances often surpass those of hyperparameter optimization [7]. Traditional techniques have relied on outlier detection methods, noise removal techniques, or ensembles of different algorithms to eliminate mislabeled instances [8]. Another approach involves removing instances that are mostly misclassified by a set of classifiers [5, 7]. Furthermore, some studies indicate that noise in class labels has a greater negative impact on the performance of the induced model than noise in input attributes [9, 10]. All of these studies indicate that the removal of detrimental instances significantly improves the quality of the learned model.
In recent years, more theoretically grounded techniques have been proposed to address label noise in large-scale datasets [11, 12]. For instance, influence functions [13] estimate how much individual samples affect model predictions, enabling principled identification of harmful points. Co-teaching strategies [14] train two networks simultaneously, each filtering out potentially noisy instances based on peer loss. Other approaches rely on training dynamics such as forgetting events [15], confidence variability [16], or margin tracking [17] to identify mislabeled data. These approaches are particularly well-suited for large-scale datasets, as they exploit patterns in model behavior that are indicative of label noise, making them practical and effective in real-world settings.
However, most of these existing approaches do not directly optimize validation loss. This is despite the fact that validation loss serves as a practical and accessible proxy for the generalization loss, which is the ultimate goal in supervised learning. If the validation loss could be directly optimized, the training subset would be refined in a more comprehensive way. It would not only filter out mislabeled instances but also address class imbalance, eliminate samples with unreliable features, and mitigate other factors detrimental to generalization. Direct optimization of validation loss is theoretically principled and practically appealing. However, this approach is seldom explored in the field because identifying the optimal training subset with respect to validation loss amounts to solving an inherently difficult combinatorial optimization problem.
In recent years, surrogate model-based black-box optimization (BBO) [18, 19, 20, 21] has gained attention as an efficient strategy for searching combinatorial spaces where the objective function is not explicitly known. The surrogate model constructs a quadratic approximation of the unknown objective function based on previously evaluated points. This approximation enables efficient selection of the next candidate solution in the combinatorial space. In particular, surrogate model-based BBO has been shown to work effectively in combination with annealing-based samplers [18], especially well with quantum annealing samplers [19, 20, 21] running on physical quantum annealers like D-Wave Advantage [22]. Quantum annealing [23, 24, 25] has traditionally been applied to solve combinatorial optimization problems [26, 27], but it is also well suited for identifying high-quality candidate points due to its demonstrated capability to sample diverse and near-optimal solutions efficiently [28, 29]. However, this method has not yet been explored in the context of the dataset cleansing or noisy-label removal. Although the number of qubits available on the current quantum annealers remains limited to the order of thousands as of 2025 and not ready to handle real-world datasets yet, this number is steadily increasing. This study aims to provide preliminary insights into the future applicability of quantum-assisted optimization for data quality improvement, using a well-controlled experimental setup.
In this work, we apply the surrogate model-based BBO with postprocessing [30] for cleansing contaminated training datasets, and evaluated its effectiveness using a well-controlled experiment called a noisy majority bit task. In the postprocessing phase, previously explored solutions are discarded from the set of sampled candidates. If a high-quality, unexplored solution is available, it is selected as the next candidate. Otherwise, a random sample is chosen to encourage further exploration. The performance of our approach is assessed using three different annealing based-samplers: the simulated annealing sampler, the simulated quantum annealing sampler, and quantum annealing sampler operating on D-Wave’s actual quantum hardware. We hypothesize that the diversity of samples generated by the quantum annealer yields a broader set of near-optimal solutions, including those not previously explored, thereby outperforming classical simulators in this context. Our approach utilizes (1) a validation loss metric for selecting data points under the assumption of a clean validation dataset, (2) a sequential training subset selection strategy utilizing BBO with a surrogate model estimating validation loss, and (3) the unique capabilities of quantum annealers to sample diverse and high-quality solutions efficiently. Together, these components provide a principled and effective framework for addressing the challenge of noise-label removal.
@@ Methods
To evaluate the effectiveness of the proposed algorithm in removing mislabeled instances, we designed a classification task called the noisy majority bit task. This task aims to predict the majority bit in a given binary input vector. The dataset for this task is partitioned into distinct, non-overlapping training, validation, and test sets to ensure an unbiased evaluation. For instance, the correct target label is 0 for the binary input vector $[0,0,0,1,0,1,0,0,1]$, whereas $1$ for $[1,1,0,1,0,1,1,0,0]$. Under ideal conditions, where all training instances are accurately labeled according to the majority bit, this task would be straightforward. However, the actual training set includes noisy instances where the labels are flipped to the minority bit, introducing mislabeled data. This contamination makes the noisy majority bit task non-trivial, providing a controlled and challenging framework for evaluating the proposed algorithm’s effectiveness in noise removal.
@@
Although validation loss can be used as a proxy for the generalization performance of a trained model, it is computationally prohibitive to evaluate the validation loss for all possible subsets of the original training dataset. To address this issue, we employed surrogate model-based BBO [18, 19, 20, 21] with postprocessing [30] to efficiently identify training subsets with low validation error in an iterative manner. Although surrogate model-based BBO with postprocessing is not itself novel, its application to noise removal—combined with the use of diverse samples from a quantum annealer during the acceptance process—is a key contribution of this work. The following part explains the detailed procedure of this algorithm and its application to the removal of mislabeled training instances. The pseudo code of the algorithm is shown in Algorithm 1.
@@
[…]
3:for $k=1$ to NN do
@@
to estimate the validation loss of the model trained by the filtered training subset $D_{\mathrm{train}}(\bm{q})$. In the surrogate model-based BBO, the surrogate model plays two roles: one is to approximate the unknown cost function with a quadratic binary form and the other is to generate binary samples with sufficiently low cost values. Here, the surrogate model $f_{\tilde{\bm{\alpha}}}(\bm{q})$ is parameterized by the parameter vector $\tilde{\bm{\alpha}}\in\mathbb{R}^{p}$ given in the form of Eq. (4) where p=1+n+n(n1)/2p=1+n+n(n-1)/2. The elements in this parameter vector $\tilde{\bm{\alpha}}$ are also present in the constant-shifted quadratic unconstrained binary optimization (QUBO) form in Eq. (2). Eq. (2) is composed of the first term $\alpha_{0}$ which represents a constant shift and the second quadratic term which represents linear interaction by its diagonal elements and pair-wise interaction by its off-diagonal elements. In short, the unknown true cost function $f(\bm{q})$ over $\bm{q}\in\{0,1\}^{n}$ is approximated by the second order Taylor expansion around $\bm{q}$, which is $f_{\tilde{\bm{\alpha}}}(\bm{q})$ in the form shown in Eq. (3). The quality of this approximation can be improved by adjusting the parameter vector $\tilde{\bm{\alpha}}$.
@@
In the first phase, we randomly sample the $n$-dimensional binary pattern $\bm{q}_{k}\in\{0,1\}^{n}$ sequentially from $k=1$ to $k=N_{0}$, and then for each binary pattern $\bm{q}_{k}$, we compute the validation loss $L_{\mathrm{valid}}(\bm{q}_{k})$ of the model trained by the filtered training subsets $D_{\mathrm{train}}(\bm{q}_{k})$. The validation loss $L_{\mathrm{valid}}(\bm{q}_{k})$ might not be distributed normally, and it is often preferable to transform it by a monotonically increasing function $g$ such as the log function to make it more normally distributed when this value becomes the target of some model and this is the case in our algorithm. Let us define the transformed validation loss as $l_{k}=g(L_{\mathrm{valid}}(\bm{q}_{k}))$. If we define $\tilde{\bm{q}}$ as in Eq. (4), the transformed validation loss $l_{k}$ can be approximated by the surrogate model $f_{\tilde{\bm{\alpha}}}(\bm{q}_{k})$ in the form of Eq. (5), which is linear in $\tilde{\bm{q}}$. Therefore, if we collect the transformed validation losses $\{l_{1},\ldots,l_{N_{0}}\}$ and the corresponding binary patterns $\{\bm{q}_{1},\ldots,\bm{q}_{N_{0}}\}$, we can update the parameter vector $\tilde{\bm{\alpha}}$ to minimize the transformed validation losses $\{l_{1},\ldots,l_{N_{0}}\}$. This optimization can be done in many forms such as the Bayesian linear regression with sparse prior [18, 31, 32], the LASSO regression [33], and the Ridge regression to update the parameter vector $\tilde{\bm{\alpha}}$ depending on the sparsity assumption of the given problem. This first step is necessary to obtain a good initial guess of the parameter vector $\tilde{\bm{\alpha}}$.
@@
In the second phase, we sequentially accept one good sample per step and compute the parameter vector $\tilde{\bm{\alpha}}_{k}$ with sequentially updated input-output pairs $\{(\tilde{\bm{q}}_{j},l_{j})\}_{j=1}^{k-1}$ available at each time step $k\in\{N_{0}+1,N_{0}+2,\ldots,N\}$. The good sample is selected as follows. After drawing $M$ samples $\{\bm{q}_{k}^{(m)}\}_{m=1}^{M}$ from a sampler at step $k$, the one with the lowest energy that is not included in the accepted sample set $S_{k-1}\equiv\{\bm{q}_{1},\ldots,\bm{q}_{k-1}\}$ is selected. If no such sample exists, a random sampling is conducted until a previously unseen sample is found, which is then accepted. The accepted sample $\bm{q}_{k}$ at step $k$ is then used to compute the transformed validation loss $l_{k}$ at step $k$, by transforming the validation loss $L_{\mathrm{valid}}(\bm{q}_{k})$ of the model trained by a filtered training subset $D_{\mathrm{train}}(\bm{q}_{k})$. Newly obtained quantities at time step $k$ such as $\bm{q}_{k}\in\{0,1\}^{n}$, $\tilde{\bm{q}}_{k}\in\{0,1\}^{p}$, $l_{k}\in\mathbb{R}$ are stored for the next iteration. At the end of the second phase, the best sample $q_{k^{*}}$ with the lowest transformed validation cost is reported where $k^{*}=\mathrm{argmin}_{k}l_{k}$.
@@
In our experiment, we used the following parameters and settings. $n=128$, $N_{0}=64$, $N=320(=64+256)$, $M=512$, $p=8257=1+128+128\times(128-1)/2$. We used a logistic regression model as a task-specific base model for solving the noisy majority bit task. The reason for using this model is that it is simple, interpretable, and fast to train. As a linear model for estimating $\tilde{\bm{\alpha}}$, we used the ridge regression model instead of the Bayesian model with sparsity prior or the Lasso model, which is usually used in the context of the surrogate model-based BBO to impose the sparsity constraint on the QUBO matrix. The reason for not using the sparse version of the model is that the data is generated in such a way that half of the training data points have noisy labels, and we expect the resulting QUBO matrix can be well approximated without sparsity assumption.
Complexity analysis
The computational cost of the proposed algorithm mainly arises from training the surrogate model and executing the sampling loop. The surrogate model involves $p=O(n^{2})$ parameters for a training dataset of size $n$. At each optimization step $k$ with $N_{0}<k\leq N$, the regression update uses all k1k-1 previously evaluated samples, which requires operations on a $(k\times p)$ data matrix. In the worst case, solving ridge regression from scratch at step $k$ costs $O(kp^{2}+p^{3})$. Summed over $N$ optimization steps, this yields a total worst-case complexity of $O(N^{2}p^{2}+Np^{3})$, or equivalently $O(N^{2}n^{4}+Nn^{6})$.
The sampling loop contributes an additional $O(NMSn)$ operations, where $M$ is the number of reads (samples) drawn per step and $S$ is the number of sweeps per read. For classical samplers (Neal SA or OpenJij SQA), one read with a fixed $S$ visits all $n$ binary variables per sweep with incremental energy updates, yielding a per-read cost of $O(Sn)$ and a total sampling cost of $O(NMSn)$. In contrast, for the D-Wave QA clique sampler, the per-step wall-clock time (including embedding) is nearly constant, so its sampling cost scales as $O(N)$. These asymptotic bounds are consistent with the measured mean per-step times in Fig. 9 ($1.196$ s for OpenJij SQA, $0.104$ s for Neal SA, and $0.011$ s for D-Wave QA clique), which show that sampling dominates the runtime for classical samplers under our settings, whereas the quantum sampler contributes only a small sampling overhead.
Although substituting $(n,N,M,S)=(128,320,512,1000)$ into the asymptotic forms may suggest that the surrogate term $O(N^{2}n^{4}+Nn^{6})$ dominates $O(NMSn)$, wall-clock runtime is governed by constant factors and implementation efficiency. In our setting, ridge regression is performed using the highly optimized scikit-learn implementation, which results in small constant factors. By contrast, SA/SQA must generate $M$ reads with $S$ sweeps per read, incurring substantial per-read overhead. Consequently, the sampling loop is the primary bottleneck for classical samplers, as shown in Fig. 9.
@@ Results
Figure 2: Instance removal patterns acquired by three different samplers over 32 runs. Each cell in the grid represents the inclusion or exclusion status of a training instance for a specific run. For any sampler, a white cell at the $r$-th row and $i$-th column indicates the $i$-th training instance is included in the filtered training set and used for actual training in the $r$-th run. The black cell, on the other hand, shows the same training instance is excluded from the filtered training set and not used for training in the associated run. The red dotted line marks the border between clean and mislabeled training instances. Seed values are changed from 0 to 31. This consistent filtering pattern provides clear evidence that the proposed algorithm can reliably identify mislabeled instances while preserving clean instances.
@@
A logistic regression model trained on the cleansed training dataset consistently demonstrated improved performance across all 32 runs as shown in Fig. 3. Here, we focus on analyzing the results obtained using the D-Wave QA clique sampler, as similar conclusions are likely applicable to the other samplers. The training errors computed on the optimized training subsets were consistently lower than the training errors computed on the original training sets without noise removal. This indicates that the mislabeled instances were successfully filtered out from the contaminated training dataset. Regardless of successful removal of the mislabeled instances, the training errors computed on the filtered training subsets remained consistently higher than the validation and test errors. This discrepancy arises because some mislabeled instances persist in the filtered dataset even after noise removal by the proposed algorithm. Please note that the validation error is lowest because it is used as the objective function of the surrogate model-based BBO, and all instances in the validation dataset are noise free. The low test errors suggest that the proposed algorithm effectively removed mislabeled instances in a way to guarantee the generalization of the trained model.
@@
Figure 3: Performance of the trained model on four different datasets. In each run, a model was trained using a filtered training dataset and then evaluated on four different datasets. The model performance was measured in log-loss error metric. Each gray dot represents one of the 32 log-loss errors for a specific dataset. Variations in error values within the same dataset are due to differences in the training instances selected across different runs by our dataset optimization algorithm. The labels Train (all) and Train (optimized) indicate whether all or filtered instances in the training dataset were used for model evaluation, respectively. The whiskers extend to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot. The two box plots on the left confirm that training errors are consistently reduced after optimizing the training dataset in all 32 runs. The two box plots on the right further indicate that test errors remain low and comparable to directly optimized validation errors, offering strong evidence that training with the optimized dataset enhances generalization performance.
@@
Further analysis revealed that mislabeled instances with a greater negative impact if not removed are more likely to be eliminated in an input-feature-dependent fashion. Fig. 2 shows that the difference in removal probabilities of fake instances is persistent over different runs and different samplers. In other words, there is a pattern in removal probability of fake instances. Fig. 4 shows the 64 input vectors contained in the fake training instances and their characteristics, such as the summed value of each input vector, together with their removal probabilities computed after 32 runs with the D-Wave QA clique sampler. The relationship between the summed values of input patterns and the removal probabilities of their associated instances is shown in Fig. 5(a). Specifically, the summed input pattern, $s_{i}=\sum_{c=1}^{9}x_{i,c}$, or equivalently the probability of an element of the input vector taking the value of 1, $p_{i}=s_{i}/9$, is shown to influence the removal probability of the input pattern $\bm{x}_{i}\in\{0,1\}^{9}$. To clarify this relationship, Fig 5(b) presents the absolute deviance of summed input pattern from the optimal threshold of 4.5, which is optimal for solving the majority bit task. This absolute deviance defined as $d(\bm{x}_{i})=|9p_{i}-4.5|$ is computed for 64 unique input patterns in the training set $\{\bm{x}_{i}\}_{i=1}^{64}$. Notably, this absolute deviance is inversely related to the entropy of a binary input pattern, defined as $s(\bm{x}_{i})=-p_{i}\log p_{i}-(1-p_{i})\log(1-p_{i})$. Therefore, Fig 5(b) demonstrates that mislabeled input patterns with lower entropy, which correspond to higher absolute deviance, are more likely to be removed. This trend may be attributed to the stronger negative impact of low-entropy patterns with incorrect labels (e.g., $[0,0,0,0,0,0,0,0,0]$ labeled as $1$) on the trained model, compared to high-entropy patterns with incorrect labels (e.g., $[0,1,0,1,0,1,0,1,0]$ labeled as $1$).
@@
Figure 4: Characteristics of the input patterns of the fake training instances and their removal probabilities using the D-Wave QA clique sampler. 1) The 9 dimensional binary input patterns of the fake training instances. 2) The summed values of input patterns, with the horizontal dashed red line marking the optimal decision boundary of 4.5. 3) The absolute deviance (or distance) of summed input patterns from the optimal decision boundary of 4.5. 4) The removal probabilities of incorrectly labeled fake instances over 32 runs with different seed values ranging from 0 to 31. 5) The removal patterns of fake training instances across 32 different runs. This removal pattern is also depicted on the right side of the dashed red line in the D-Wave QA clique result shown in Fig. 2. The relationship between characteristics of input patterns and their removal probabilities is further analyzed in Fig. 5, highlighting how input features influences the likelihood of removal.
@@
Figure 5: Relationship between the removal probabilities and the characteristics of 64 input patterns of the fake training subset. The removal probabilities are compared with (a) the number of ones in each input pattern of the fake training subset and (b) the absolute deviance from the optimal threshold of 4.5 to the summed values of input patterns. In both plots, the whiskers are drawn to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot. Both figures clearly illustrate that mislabeled training instances with greater detrimental impact on the trained model—for example, a zero vector mislabeled as class 1—are more likely to be removed.
@@
Figure 6: Comparison of the three samplers in terms of surrogate losses and the types of sequentially obtained samples. Each surrogate loss models the log-transformed validation loss, which takes continuous values in the real domain. The first 64 steps up to the dotted gray line are called the initialization phase, during which samples are obtained randomly. The remaining 256 steps constitute the optimization phase, where samples are obtained using the surrogate model-based BBO with postprocessing. In each optimization step, 512 samples are drawn from a selected sampler. These samples are sequentially checked from the one with the lowest energy to the one with the highest energy to see if they are already contained in the accepted sample set. The first sample not contained in the accepted sample set is accepted and inserted into the accepted sample set. If all 512 samples are already contained in the accepted sample set, a new sample not contained in the acquired sample set is drawn randomly. The trend of the lowest energy up to that step is shown with a blue solid line. A sample type is marked as optimal if the accepted sample has the lowest energy among the 512 samples. A sample type is marked as suboptimal if it is from the 512 samples but does not have the lowest energy. A sample type is marked as random if it is not obtained from the 512 samples and is drawn randomly. In all three cases, all parameters are fixed except for the solver types. As shown in Fig 6(c), the D-Wave QA clique sampler effectively exploits diverse samples obtained from a physical quantum annealer, thereby achieving superior solution quality compared with classical samplers such as OpenJij SQA (Fig 6(a)) and Neal SA (Fig 6(b)) samplers.